Allen is 4 yr older than his brother Randy. Three years ago, Allen was twice as old as Randy. How old is each now? Please solve using system of equations if possible please.
Answer by lwsshak3(11628)

Allen is 4 yr older than his brother Randy. Three years ago, Allen was twice as old as Randy. How old is each now? Please solve using system of equations if possible please.
***
let x=Allen's present age
let y=Randy's present age
x=y+4
..
3 yrs ago:
x-3=Allen's age
y-3=Randy's age
x-3=2(y-3)
...
x=y+4
x-3=2(y-3)=2y-6
..
x-y=4
x-2y=-3
subtract:
y=7
x=7+4=11
..
Allen's present age=11
Randy's present age=7